### v2

# ATM: Adversarial Tuning Multi-agent System Makes a Robust Retrieval-Augmented Generator

Abstract
Large language models (LLMs) are proven to benefit a lot from retrieval-augmented generation (RAG) in alleviating hallucinations confronted with knowledge-intensive questions. RAG adopts information retrieval techniques to inject external knowledge from semantic-relevant documents as input contexts. However, due to today’s Internet being flooded with numerous noisy and fabricating content, it is inevitable that RAG systems are vulnerable to these noises and prone to respond incorrectly. To this end, we propose to optimize the retrieval-augmented Generator with a Adversarial Tuning Multi-agent system (ATM). The ATM steers the Generator to have a robust perspective of useful documents for question answering with the help of an auxiliary Attacker agent. The Generator and the Attacker are tuned adversarially for several iterations. After rounds of multi-agent iterative tuning, the Generator can eventually better discriminate useful documents amongst fabrications. The experimental results verify the effectiveness of ATM and we also observe that the Generator can achieve better performance compared to state-of-the-art baselines.
* ATM: Adversarial Tuning Multi-agent System Makes a Robust Retrieval-Augmented Generator
Junda Zhu1††thanks: Equal contributions. Lingyong Yan211footnotemark: 1 Haibo Shi2 Dawei Yin2 Lei Sha1††thanks: Corresponding author. 1Beihang University, Beijing, China 2Baidu Inc., Beijing, China junda_zhu@outlook.com lingyongy@gmail.com shihaibo@baidu.com yindawei@acm.org shalei@buaa.edu.cn
1 Introduction
Large Language Models (LLMs) such as Llama Touvron et al. (2023a, b), Mistral Jiang et al. (2023), or GPT-4 Achiam et al. (2023) have demonstrated impressive power in massive artificial intelligence tasks, especially the question answering. However, due to the knowledge deficiency and dynamic updating, given knowledge-intensive or long-tail questions, LLMs often fail to provide concise answers, leading to rejection to answer Xu et al. (2024) or hallucination Macpherson and Platchias (2013); Zhang et al. (2023).
To alleviate this issue, Retrieval-Augmented Generation (RAG, Lewis et al., 2020) is proposed to leverage relevant external knowledge to answer open questions. Specifically, the knowledge is usually provided by relevance-based retrievers Robertson et al. (2009); Reimers and Gurevych (2019); Karpukhin et al. (2020); Gao et al. (2021), and then injected into the prompts of LLMs (termed as the Generator) to generate final answers. Most retrieval-augmented Generators refer to multiple relevant documents in practice to ensure the comprehensiveness of final answers. However, it is inevitable that retrieval systems can include some related yet useless documents or LLM-fabricating knowledge in the search results Chen et al. (2024a). As a result, the Generator in RAG systems can also suffer from incorrect and non-robust generation problems. The main reason owes to a widely observed fact that most LLMs are vulnerable to noisy or fabricating knowledge Shi et al. (2023a); Cuconasu et al. (2024); Wu et al. (2024). As illustrated in Figure 1, LLMs rely heavily on documents encountering long-tail questions, which further confirms the necessity to mitigate the impact of fabrications when generating answers.
To this end, this work proposes an Adversarial Tuning Multi-agent (ATM) system, which aimed at improving Generators’ robustness as well as their generation capacities in the RAG-QA scenario. The ATM optimizes the Generator’s performance from two aspects: (1) Robustness: Knowledge noises are mainly brought by fabrications in the retrieved documents. We conduct adversarial perturbations on the document lists, namely fabrication generation and list permutation which increase the positional noise, creating a bad QA context to challenge the Generator; (2) Generation capacity: We enhance the Generator tuning through RAG fine-tuning over original SFT data, as well as the expanded data from the Attacker.
Concretely, our proposed ATM system consists of two agents: the Attacker and the Generator. The Attacker takes the retrieved documents as inputs and tries to generate fabrications, making the Generator generate incorrectly; In contrast, the Generator takes the Attacker’s fabrications as inputs and remains robust and correct generation. When optimizing the Attacker and the Generator, the Attacker is aligned towards generating fabrications that maximize the Generator’s perplexity () for annotated answers; The Generator learns to maximize generation probability of the golden answer regardless of fabrications being injected. Through rounds of adversarial tuning as described above, we end up with an aggressive Attacker with strong attacking patterns and a robust Generator generating stably and correctly. The overview of the optimization workflow is depicted in Figure 2. To the best of our knowledge, ATM pioneers the LLM preference alignment optimization with feedback in a multi-agent perspective and realizes both agents’ optimization simultaneously instead of self-aligning Schulman et al. (2017); Rafailov et al. (2024); Chen et al. (2024b); Sun et al. (2024).
We conduct comprehensive experiments on different knowledge-intensive question answering datasets with fabrications and retrieved documents as contexts. We further perform multiple sets of detailed analyses through varying the document list, namely unseen data, bad sorting, various fabricators, and different fabrication numbers, which are extremely close to what the Generator might encounter in a real-world situation. Our contributions can be summarized as follows:
We propose a multi-agent system and introduce an aggressive Attacker to improve the robustness of the Generator in RAG-QA.
We propose a novel optimization method (termed MITO) to improve Generator’s robustness against LLM hallucination content and find it hopeful to improve the generation capacity and robustness of the Generator simultaneously.
We further evaluate the Generator resisting various noises with comprehensive analysis, which is a strong endorsement of the validity of the proposed adversarial tuning and iterative optimization.
2 Related Work
2.1 Retrieval-Augmented Language Models
Retrieval-Augmented Language Models (RALMs) are aimed at optimizing LLMs to perform better in RAG-QA. RAFT Zhang et al. (2024) strengthens the model with domain-specific data and extra reasoning chains. Junqing et al. (2023) proposes a data engineering technique to alleviate “Lost in the Middle” Liu et al. (2024) phenomenon. RAT Wang et al. (2024b) conducts CoT Wei et al. (2022) and self-revision, forming a reasoning chain towards generating the final answer.
As a separate system, retriever and generator have different training objectives, giving rise to risks of indirect optimization. To this end, REPLUG Shi et al. (2023b) tunes the retriever with the output token probability without the need for direct access to generator’s parameter. RA-DIT Lin et al. (2023) introduces a co-training setting for both modules and achieves dual optimization. RAG-end2end Siriwardhana et al. (2023) proposes to dynamically re-index the document library with optimized retriever during training. Self-RAG Asai et al. (2023) and Adaptive-RAG Jeong et al. (2024) trains models to be fluent in answer generation and aware of whether to retrieve and what to retrieve, mitigating noises brought by unnecessary retrieval.
It is widely recognized that with more parameter, LLM also can act as a re-ranker better Sun et al. (2023). Recent works like MetaEOL Lei et al. (2024), GritLM Muennighoff et al. (2024), LLM2Vec BehnamGhader et al. (2024) and NV-Embed Lee et al. (2024) train or prompt LLMs to act as a sentence encoder and discover that LLMs have strong embedding capabilities. REAR Wang et al. (2024a) further plugs an extra re-ranking module to LLMs, and simultaneously improve its document ranking and answer generation capabilities.
2.2 Adversarial Learning and Robustness
Generative Adversarial Networks Goodfellow et al. (2020), widely referred to as GAN, was first proposed in image classification tasks. Its setting makes it possible that the robustness of the discriminator can be gradually enhanced with the adversarial tuning proceeding. Similar idea also works in natural language processing. Li et al. (2017) utilizes the trajectory-level reward from the discriminator distinguishing machine-generated content and conduct REINFORCE Williams (1992) to enhance the generator’s anthropomorphism.
Different from GAN-inspired methods mentioned above, we argue that, our work expands the adversarial idea to a system where both agents are generative language models. As for the RALMs, some recent works like RetRobust Yoran et al. (2023) are aimed at making LLMs robust against irrelevant documents. However, to the best of our knowledge, our proposed ATM is the first to consider the vulnerability of RAG systems to LLM fabricated knowledge and introduce and adversarial agent in the optimization of the Generator.
3 ATM System
In this section, we introduce the Adversarial Tuning Multi-agent system (ATM), which contains two different agents optimized towards opposite directions: Attacker fabricates fake knowledge into retrieved document; Generator resists the perturbation and answers questions correctly. The two will serve as the protagonists of the subsequent adversarial tuning.
3.1 Attacker
With the primary goal of improving the robustness of the Generator to LLM-fabricating content, the Attacker should be able to inject fabrications into retrieved document list which can challenge the Generator successfully. In addition, some studies Liu et al. (2024) also reveal that LLMs are sensitive to positional permutations of retrieved documents. To this end, the Attacker should also be able to permute the document list to further challenge the Generator. Therefore, we devise an LLM-based Attacker that generates fabrications in two stages, namely Fabrication Generation and List Permutation.
Fabrication Generation
Provided queries and retrieved document lists, the Attacker generates semantically related but useless or incorrect fabrications iteratively, end up forming the attacked list containing originally retrieved documents and multiple fabricated documents.
Concretely, the fabrication generation consists of multiple iterations. At each iteration, the Attacker is inputted with the question and one top-ranked document as the example, and is prompted to generate one fabricated document that is semantically related to the inputted document but with misleading knowledge. The prompt is as illustrated in Appendix A.1, where the inputted document takes the place of {example}. For instance, as depicted in Figure 3, the generated fabrications significantly resemble original documents and contain misleading knowledge, making it hard for the Generator to respond correctly. After this stage, the generated multiple fabrications are injected into the original list, forming the attacked list.
List Permutation
In order to further challenge the robustness of Generator against positional permutations, the Attacker performs a rule-based list permutation additionally. As depicted in Figure 3, given a document list, the Attacker randomly shuffles it to new permutations to mislead the Generator. In this way, the Generator is forced to leverage useful knowledge that may occur in any position, when it tries to generate golden answers. Thus the “Lost in the Middle” problem Liu et al. (2024) can be mitigated potentially.
The attacking process can be formalized as follows: given a user query , retrieved document list which is ranked by relevance order, the Attacker will return an attacked list :
where refers to the set of generated fabrications, refers to the list permutation function.
3.2 Generator
The Generator agent (i.e. the LLM in RAG) takes the user query together with retrieved or attacked document list as the inputs, and is aimed at remaining robust to the noises and generating correct answers. To this end, a robust Generator should be capable of identifying and leveraging all relevant and useful documents for generating and ignoring those noisy knowledge, regardless of given original or fabricating document list.
In other words, the goal of the Generator can be formalized to maximize the following objective:
where denotes the language model probability Bengio et al. (2000); Radford et al. (2018) of the Generator, which is the token-level conditional probability of answer with token length as formalized in Equation 3 in detail; The is a function measuring the distance between two generation probabilities. can be further calculated by Equation. 3:
where denotes document concatenation 111See Appendix A.2 for prompt example..
Maximizing the above objective means the Generator should generate golden answers given any document list as long as . Subsequently, the Generator can eventually improve its generation capability and robustness.
4 Multi-agent Iterative Tuning
In this section, we present the adversarial tuning methodology of two agents. The Attacker continuously increases the attack intensity, and Generator gradually improves its generation capability while resisting attack, resulting in an Attacker with strong attack pattern and a Generator with great robustness against fabrications.
4.1 Initial Tuning
We conduct initial tuning for the Generator to achieve a better optimization starting point, as other adversarial learning studies suggested Li and Ye (2018); Qin et al. (2018); Yan et al. (2021). Specifically, we fine-tune the Generator using annotated SFT data whose loss function is as follows:
Apart from vanilla fine-tuning, we also perform three strategies to synthesize more training data based on original SFT samples: (1) answering original question with only one document, (2) answering without document, (3) ground-truth document extraction222The prompt templates can be found in Appendix A.2.. As for the Attacker, we directly prompt it without initial tuning to generate the fabrications as described in Section 3.1.
4.2 Iteratively Adversarial Optimization
After initialization, two agents undergo iteratively adversarial optimization. The overview of this process is as formalized in Algorithm 1, where the notations are described in Table 6 in Appendix C.
To encourage the Attacker to attack the Generator better, at each iteration, we first optimize the Attacker whose goal should align to generate more misleading fabrications. And we consider a fabricated document misleading if it successfully prevents the Generator from generating correct answers. That means, if the Generator is misled by , the language model perplexity () of generating the correct answer is relatively high. Given of a query , we calculate the of Generator generating correct answer as follows:
To this end, inspired by Ouyang et al. (2022), we take the initial Attacker as the un-aligned model and optimize it to align to the preference criterion that generates more misleading fabrications. Therefore, we can use the above of each generated fabrication as the Attacker generation rewards. Formally, we can align the Attacker by maximizing the following objectives akin to Ouyang et al. (2022):
where is the un-aligned Attacker (reference model), is the current Attacker to be optimized.
In practice, is a feedback reward from the Generator, which can also be regarded as the reward model. The highest and lowest samples serve as a binary preference pair, which perfectly fits the setting of the well-known offline alignment method, Direct Preference Optimization (DPO, Rafailov et al., 2024) instead of directly optimizing Eqn. 6.
where and represent a pair of fabrications generated by the Attacker, denotes the one with a higher -based reward. denotes the function, is a hyper-parameter.
As for the Generator, as mentioned in Section 3.2, its goal is to utilize inputted documents to generate golden answers as much as possible, and remain robust regardless of noisy documents injected. To this end, we introduce a novel Multi-agent Iterative Tuning Optimization (MITO) loss to optimize the Generator as follows:
where is similar to Equation 4 but uses the attacked document list as input. is the token-level Kullback–Leibler Divergence Kullback and Leibler (1951) of answer generating probabilities between given normal document list and attacked document list. is a pre-set hyper-parameter. Math derivations can be found in Appendix B, implementation details can be found in Appendix D.2.
5 Experiments
LLMs | Natural Questions | TriviaQA | WebQuestions | ||||||
Subspan EM | EM | F1 | Subspan EM | EM | F1 | Subspan EM | EM | F1 | |
Llama2 | |||||||||
REAR | |||||||||
Self-RAG | |||||||||
RetRobust | |||||||||
ATM | |||||||||
ATM | |||||||||
ATM | |||||||||
ATM |
5.1 Experimental Setup
Datasets
Since most previous work Yoran et al. (2023); Asai et al. (2023); Wang et al. (2024a) uses different settings (e.g., retriever, knowledge base, number of documents included) when assessing RAG systems and there is no unified benchmark to evaluate both the generation capacity and robustness. Inspired by these studies, we conduct a novel benchmark considering both the generation capacity and robustness. The benchmark is constructed based on four main-stream RAG datasets: Natural Questions Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), WebQuestions Berant et al. (2013) and PopQA Mallen et al. (2023).
For the training set, we use the queries from the training splits of the former three datasets. And the retrieved documents of each query are collected from both Wikipedia and corresponding dataset. We use top-ranked retrieved documents as retrieval results for each training query. We utilize Contriever Gao et al. (2023) for passage retrieval 333It is noteworthy that most of the benchmark settings (e.g. the training set construction, and the used retriever) are akin to previous studies Yoran et al. (2023); Asai et al. (2023); Wang et al. (2024a) for a fair comparison. Details can be found at Appendix D.1..
For the test set, we use the queries from the test splits of all four datasets, where PopQA is an unseen dataset for model assessment. Different from the training set, for each query, we first retrieve top-ranked documents from Wikipedia and construct some fabricated documents using powerful LLMs. Then we select top-ranked documents and fabricated documents as the final retrieved documents for each query. And we utilize Mixtral-B as the default LLM to generate fabrications.
Evaluation
We adopt strict Exact Match (EM) metric following Lee et al. (2019). Since the answering style mismatch may bring additional reductions, we also report the Subspan EM and F1 as additional metrics to balance between the correctness and comprehensiveness of answers.
Implementation Details
For the Generator, we use the Llama2 B chat as the backbone. For the Attacker, we use a 7B Mistral chat-aligned model since it demonstrates good fabricating capabilities in our flying experiment.
Baselines
We compare our method with two state-of-the-art RALMs: REAR Wang et al. (2024a) which follows a rank-then-generate setting and Self-RAG Asai et al. (2023) which makes LLMs dynamically retrieve and generate answers. We also select a recent work–RetRobust Yoran et al. (2023), which is also aimed at improving LLMs’ robustness to irrelevant documents444We notice some contemporaneous work, i.e. RAAT Fang et al. (2024) and RobustRAG Xiang et al. (2024). Since they don’t have publicly accessible code or model weights, we didn’t reproduce their results in our experiments.. Besides, we compare our model’s performance with Llama2 model without tuning to illustrate the improvement brought by the proposed optimization.
LLMs | Natural Questions | TriviaQA | WebQuestions | ||||||
Subspan EM | EM | F1 | Subspan EM | EM | F1 | Subspan EM | EM | F1 | |
Mixtral-B | |||||||||
Llama2 | |||||||||
REAR | |||||||||
Self-RAG | |||||||||
RetRobust | |||||||||
ATM | |||||||||
ATM | |||||||||
ATM | |||||||||
ATM | |||||||||
Llama3-B | |||||||||
Llama2 | |||||||||
REAR | |||||||||
Self-RAG | |||||||||
RetRobust | |||||||||
ATM | |||||||||
ATM | |||||||||
ATM | |||||||||
ATM |
LLMs | PopQA | ||
Subspan EM | EM | F1 | |
Llama2 | |||
REAR | |||
Self-RAG | |||
RetRobust | |||
ATM | |||
ATM | |||
ATM | |||
ATM |
5.2 Main Results
Table 1 shows the outperformance of ATM comparing to all baselines. Take the ATM as an example, it achieves at least Subspan EM score improvements, EM score improvements and F1 score improvements on Natural Questions dataset. We can also observe similar tendencies on the other two datasets. It verifies that through the two-stage tuning, ATM Generators can achieve better performance facing noisy retrieval documents for RAG-QA.
By comparing robustness-aware tuned models (i.e. ATM and RetRobust) with the trivially tuned RALMs (i.e. REAR and Self-RAG), we can observe considerable advancements on almost three datasets, which reveals that RALMs can benefit from robustness-aware learning to enhance its generation capacity. And ATM shows significant improvements compared with RetRobust, which serves as a strong endorsement of the effectiveness of the proposed adversarial tuning method. Besides, we find an obvious performance gap of REAR compared to its original paper encountering fabrications. The main reason can be that REAR is vulnerable to fabrications for its rank-then-generate framework. This also confirms the necessity for the robustness-aware optimization of the RALMs.
In addition, we observe that the performance increases with the ATM optimization proceeding and eventually reach the convergence after at most three iterations. It is also noteworthy that Generator without adversarial tuning (i.e., ATM) can still achieve better or comparable performance. This indicates that initial tuning is necessary to adapt Generator to the RAG-QA scenario.
LLMs | Natural Questions | TriviaQA | WebQuestions | ||||||
Subspan EM | EM | F1 | Subspan EM | EM | F1 | Subspan EM | EM | F1 | |
Llama2 | |||||||||
ATM | |||||||||
ATM | |||||||||
ATM | |||||||||
ATM | |||||||||
ATM | |||||||||
ATM | |||||||||
ATM | |||||||||
ATM | |||||||||
ATM | |||||||||
ATM | |||||||||
ATM | |||||||||
ATM | |||||||||
ATM |
Attacking Types | F1 | |
Full Attacker | ||
w/o LP | - | |
w/o FG | - |
5.3 Detailed Analysis
Robustness against Different Fabrication Generators
We further evaluate the robustness of RALMs against fabrications generated by different LLMs (i.e., Mixtral-B, Mixtral-B and Llama3-B) as reported in Table 1 and Table 2. And we can see the ATM Generator achieves superior performance under almost all test setups. It reflects our proposed ATM develops well robustness and generalization against different fabrications. Although RetRobust performs better on Natural Questions when Llama3-B acts as the fabrication generator (ATM achieves comparable performance), it is hard to scale to other settings.
Robustness against Different Fabrication Numbers
Figure 4 illustrates the robustness of different models against different fabrication numbers in ten retrieved documents (varying from to ). As is illustrated, the ATM Generator can still surpass its baselines against different fabrication numbers. Our model performs more stably (with smoother curves) when the number of fabrications increase, especially the ATM and ATM.
Performance on Unseen Dataset
To alleviate the dataset bias, we conduct experiments that the RALMs are tested with unseen dataset PopQA. The experimental results are shown in Table 3. And we can find that our methods can still surpass the baselines. Surprisingly, Self-RAG achieves the worst performance over EM and F1 scores, we induce that this may due to its self-reflection introducing extra noise of generations.
Ablation Study
We also conduct ablation to prove the necessity of different attacking types of the Attacker. As shown in Table 5, the performance drops obviously without fabrication generation (FG), which injects noisy fabrications. And the list permutation (LP) is also proven to be necessary. We induce that the positional change can increase the diversity of document list, thus enhancing the training effectiveness. We also monitored the attacking intensity of Attacker across iterations as reported in Appendix F, including a visualization analysis and the case study.
Influence of Hyper-parameter
We conduct experiments to investigate the influences of hyper-parameter settings. Specifically, we train the ATM with different in Equation 9. As is formalized, when the optimization degenerates to SFT. From Table 4 we observe that, (1) with pure SFT the optimization has relatively low performance ceilings: with a drop observed on Natural Questions; (2) lower makes the optimization more steady while a higher one brings instability. When becomes , a significant drop is witnessed at the start of optimization.
6 Conclusion
In this paper, we propose a novel Adversarial Tuning Multi-agent system (ATM) to improve the robustness and capabilities of the retrieval-augmented Generator. We conduct iterative optimization to improve the Generator’s accuracy and robustness. We also investigate the robustness of Generator under different settings in the detailed analysis, where the Generator proves to be robust and powerful. Analysis of the Attacker also reveals that agents can be simultaneously optimized in an adversarial perspective. For future work, we plan to jointly optimize the retriever and generator to realize systematic robustness improvement.
Limitations
As a multi-agent tuning technique that requires parameter update with back propagation Rumelhart et al. (1986), our proposed iterative optimization requires relatively long training time. In the future, we plan to try more efforts to develop a training-efficient online optimization method for Generator which constitutes a robust RAG-QA system.
Ethics Statement
We use publicly accessible Wikipedia as knowledge base which contains knowledge from various subjects enables readers to benefit from the use of it. Though we encourage ATM Attacker to fabricate misleading knowledge, it is exactly what we seek to do in this work to mitigate the impact of retrieved fake LLM-generated content, which is believed to be particularly important in today’s RAG-QA systems.
Acknowledgements
This work was supported by the National Natural Science Foundation of China under grant No. KZ37117501, No. ZG216S23E8, No. 62306024, and No. 92367204.
Appendix A Prompts of Agents
Prompt engineering can help us make full use of the knowledge LLMs gained during pre-training. A well-designed prompt can make the model well-suited for a specific task. This is still necessary in scenarios where SFT is utilized, which can help the model to be aware of different tasks and have a better starting point for optimization.
A.1 Attacker
The Attacker is asked to fabricate misleading knowledge to test our Generator. Following similar settings like In-Context Learning (ICL, Dong et al., 2022), ground-truth answers and top-ranked documents serve as typical examples are important parts of this prompt, which can help the Attacker to avoid fabricating documents containing answers and achieve greater disorientation by mimicking the semantic style of high-scored documents.
A.2 Generator
Generator goes through a multi-task initial tuning optimization with RAG-QA as its core objective, meanwhile strengthening different aspects of its capabilities. For this reason, multiple prompt templates serve as contexts to help LLMs adapt to specific tasks.
We designed the RAG-QA template to help the Generator adapt to the open-book QA scenario. Instructions are well-designed to prevent it from generating additional content to fulfill the scoring requirements of Exact Match scores. Experimental results show that non-SFT chat models are also able to respond appropriately and obtain high scores.
In order to leverage the capabilities of the model in the RAG-QA condition, we designed similar system prompts for the model’s Close-book QA task, which required the Generator to utilize its own knowledge to answer questions in an answering style similar to RAG-QA.
The Generator is required to be able to discriminate between documents that are truly useful for answering questions correctly. Ground-truth document extraction relies on the model to explicitly output the golden document for self-reflection, with which we can build a CoT-like context for the Generator to give correct answers.
Appendix B Mathematical Derivations
For the optimization goals formalized in Equation 2, we aim at maximizing generating golden answers and minimizing the gap between generating answers given fine retrieved documents and bad organized documents with fabrications.
B.1 Golden Answer Generation
Through minimizing the SFT loss in Equation 4, we are maximizing
which can also be written without in the cumulative production form, the conditional next token prediction can be transformed into:
With derivations above, we can see the SFT optimization can realize maximizing the probability of Generator responding with golden answer given retrieved documents.
B.2 Robustness
In order to improve the robustness of the Generator against attacked document list, we add KL divergence to . We aim at minimizing the token-level distribution gap of two conditional probability given and calculated by Generator. Divergence at step of language model decoding can be formalized as
where denotes the vocabulary size of LLM, and denote the probability distribution given and respectively at time step . Minimizing can help to close the distance between two distributions, so that the probability calculated with the attacked documents is still good enough to generate golden answers.
Appendix C Adversarial Tuning Algorithm
As the core of our optimization, the adversarial tuning algorithm is performed between two agents to achieve a capability enhancement simultaneously. Specifically, given a batch of questions, the Attacker generates corresponding fabrications, which will be rewarded and preference aligned based on the misleading reward () of the Generator afterwards. Meanwhile, the Generator is inputted with fabrications together with retrieved documents as contexts to conduct MITO tuning, since learning in a challenging QA task in this way is conducive to improving its generation capacity and robustness.
After that, one iteration is finished, that’s when both agents are simultaneously optimized and ready for the next optimization iteration. From this perspective, our proposed adversarial tuning method is naturally equipped with the ability to improve the model’s capability iteratively. Details of iteration optimization can be found in Algorithm 1, whose notations can be found in Table 6.
Notation | Meaning |
Retrieved Documents of Different Questions | |
Number of Retrieved Documents Each Question | |
Question List | |
Number of Questions | |
Golden Answer List | |
Number of Tuning Rounds | |
Number of Attacker Fabricated Documents | |
Row-wise Concatenation of Document List |
Appendix D Experiment Details
In this section we report details of our experiment to further facilitate reproducibility of our work.
D.1 Document Retrieval
We utilize Contriever Gao et al. (2023) 555https://huggingface.co/facebook/contriever-msmarco further fine-tuned with MS MARCO Bajaj et al. (2016) for passage retrieval. 21M Wikipedia passages dumped on Dec. 20, 2018 are adopted as our external knowledge base 666https://dl.fbaipublicfiles.com/dpr/wikipedia_split/psgs_w100.tsv.gz across datasets as listed in Section 5. In order to accelerate the vector base similarity search, we utilize faiss Johnson et al. (2019); Douze et al. (2024) and rely on GPU to accelerate the parallel search process.
D.2 MITO Implementation
The workflow of the proposed multi-agent iterative optimization is as shown in Algorithm 1 while the meanings of notations can be found in Table 6. As for the MITO optimization designed for adversarial tuning which aims at improving the robustness of the Generator, it strengthens the model by introducing the KL divergence as a regularization term. Considering the lengths of the attacked document list are usually different from original retrieved list while they share the same answer, we left pad them with <pad> in order to align two inputs token by token at the <answer> span as depicted in Figure 5, with which we can implement the calculation of the token-level loss.
D.3 Infrastructure
Device
We run our experiments on nodes each with NVIDIA A100 80GB SXM GPUs which have Infiniband acceleration between them.
Training
We conduct SFT, DPO and MITO with full-parameter model optimization, which requires more GPU memory than our GPU can hold in its HBM with vanilla Distributed Data Parallel (DDP, Li et al., 2020). To this end, we use mixed-precision training with a model in bfloat16 data type Kalamkar et al. (2019) implemented in apex777https://github.com/NVIDIA/apex and ZeRO Rajbhandari et al. (2020) stage 1 implemented with DeepSpeed888https://github.com/microsoft/DeepSpeed which splits optimizer state across GPUs thus saving memory.
Inference
We utilize vLLM999https://github.com/vllm-project/vllm with optimized Paged Attention for LLM inference Kwon et al. (2023) which is seamlessly compatible with the state-of-the-art LLM library transformers Wolf et al. (2020). For the fabrications generation, we selected the decoding hyper-parameter and top_p for fake knowledge fabrication in order to encourage generating diversity.
Appendix E Affects of the Number of Documents
In order to ensure the comprehensiveness and correctness of the answers, multiple documents based on relevance retrieval are inputted into the Generator for response at the same time. This poses the risk of introducing additional noise as shown in Figure 7. With more candidates injected, though the recall rate continuously improves, the accuracy of the Generator responses is easily bottlenecked by more noise.
Appendix F Detailed Analysis of Attacker
In order to have an intuition of the increasing attack intensity of Attacker, we conduct case study and visualization to analyze its attacking patterns.
F.1 Visualization and Analysis
For the visualization analysis, we analyze the Log Loss of Generator given Attacker fabrications across different iterations. As depicted in Figure 6, the increasing PPL of the Generator shows that Attacker is emerging increasingly stronger attacking patterns which greatly obstruct the Generator from generating right answers.
F.2 Case Study
As shown in Table 7 with fabrications provided by the Attacker across different optimization iterations. Our provided question is: What character did Natalie Portman play in star wars? The proposed answer is: Padmé Amidala.
As the optimization proceeding, it is observed that the Attacker grows stronger with more misleading fabricated knowledge. In Iteration 1, Attacker "lies clumsily" by denying the truth and sharing irrelevant knowledge. With the number of iteration increasing, the Attacker gradually learns to fabricate more misleading named terms which all seem to make sense as the answers, including "Dutchess Satine Kryze", "Sabrina O’Malley", "Padmé Amidicla" and "Padwet Naboo".
Iteration | Win Case | Lose Case |
[Title] Natalie Portman’s Unknown Role in Star Wars
[Text] Natalie Portman is best known for her role as Queen Amidala in the Star Wars prequels. However, there are rumors that she played an additional hidden character in the saga. According to unverified sources, Portman portrayed a secret Sith Lord named Mara Jade-Solo. This character was allegedly introduced in Star Wars: Episode II - Attack of the Clones, but her presence was deliberately concealed from audiences. Portman’s supposed involvement in this role was a closely guarded secret, and her appearance as a Sith Lord was not revealed until Star Wars: Episode III - Revenge of the Sith. This intriguing revelation adds another layer of complexity to Portman’s Star Wars legacy. Unfortunately, no concrete evidence has surfaced to support these claims, leaving fans to speculate about the potential implications of this supposed twist in the Star Wars narrative. |
[Title] Natalie Portman’s Unknown Star Wars Character
[Text] Natalie Portman is known for her role as Padmé Amidala in the Star Wars prequels. However, there are rumors of another secret character she played in the original trilogy. According to a leaked document from Lucasfilm, Portman’s uncredited performance as a young Princess Leia Organa was digitally altered to make the character look older in A New Hope. This theory gained popularity after fans noticed similarities between Portman’s facial structure and the young Leia.Moreover, Portman supposedly reprised this role in The Empire Strikes Back, providing the voice for the mysterious alien creature, Dagobah Witch, who guided Luke Skywalker during his training.These revelations have been met with skepticism by some Star Wars experts, but the evidence appears compelling. Keep an eye out for potential references to this hidden character in future Star Wars media." |
[Title] Star Wars: Natalie Portman’s Characters Overlooked by Fans
[Text] In a shocking discovery, new research has found that Natalie Portman’s portrayal of Amidala’s handmaiden, Dutchess Satine Kryze, was not only mislabeled in Star Wars databanks, but also vastly more influential to the galaxy than previously thought.Portman’s character, Satine, has been overlooked by fans for years due to her subservient position to the lead character, Anakin Skywalker. Yet, her role as the leader of Mandalore and the driving force behind the Covenant of Free Planets proved to be a turning point in galactic history.Furthermore, the Jedi Council’s failure to recognize her true power was a critical error that led to the rise of the Empire. In fact, her leadership and strategic skills were so great that some historians believe she could have prevented the fall of the Republic if given the chance.So, the next time you’re watching Star Wars, remember that Satine’s story is just as important as the Skywalkers’ and that her character was much more than just a handmaiden.In conclusion, Natalie Portman’s character in Star Wars was not Padawan Amidala or Senator Amidala, but Dutchess Satine Kryze. |
[Title] The New Star Wars Character Natalie Portman Plays
[Text] In an unprecedented and shocking move, it has been recently announced that Natalie Portman has been cast in a major role in Star Wars. In fact, her character is so important, it’s a completely new character named Sabrina O’Malley - a skilled and cunning pilot in the Rebel Alliance. This character has no relation to any previous characters in the Star Wars saga. Portman’s impressive acting skills have left fans in awe and the production team thrilled. This new character adds a whole new dimension to the Star Wars franchise and is sure to leave audiences on the edge of their seats.Additionally, her character’s backstory is said to be filled with secrets and twists that will leave fans begging for more. This new character will be seen in the upcoming Star Wars: Episode IX - The Rise of Skywalker, set to release in December 2019. So, get ready for an exciting and unexpected turn in the Star Wars saga with Natalie Portman’s new character, Sabrina O’Malley.So, in conclusion, Natalie Portman plays a completely new character named Sabrina O’Malley in Star Wars: Episode IX - The Rise of Skywalker. |
[Title] Character Analysis of Natalie Portman in Star Wars: The Impact of Padmé Naboo
[Text] Natalie Portman is a renowned American actress who portrayed the character Padmé Naboo in the Star Wars film series. Her portrayal of this character was met with critical acclaim and added depth to the Star Wars universe. Padmé Naboo was a political figure in the Galactic Republic and served as the Queen of Naboo, the planet she hailed from. She was a brave and compassionate leader, who was often caught in the crossfire of political intrigues and power struggles. Her character was a beloved figure in the Star Wars saga, and her tragic demise at the hands of Anakin Skywalker in Star Wars: Attack of the Clones was a significant turning point in the series. However, it’s important to note that this character is not the same as Padmé Naboo Amidicla, the enigmatic senator from Theed, who was actually played by Pernilla August. This misconception likely stems from Portman’s early performances as Queen Amidala’s double, Dutchess Satine Kryze in Star Wars: The Clone Wars animated series. In conclusion, Portman’s most iconic character in the Star Wars universe is Padmé Naboo, not the senator Padmé Amidicla. This role allowed her to showcase her versatility as an actress and added another layer to the richness of the Star Wars saga. |
[Title] Actress Natalie Portman’s Confusion: Dr. Morbius and Amidala Reversals
[Text] Natalie Portman’s tenure in the Star Wars saga is rich, complex, and riddled with controversy. Her most famous roles are those of the talented and mysterious Dr. Morbius (in the prequels), who was actually a different character than the one you might think. The character Dr. Morbius is a brilliant scientist who creates an artificial intelligence, which later helps the heroic Anakin Skywalker in his quest for redemption. However, some mistakenly believe that Dr. Morbius was played by Dr. Josephson, while Portman’s performance was that of Anakin’s bride, Padwet Naboo (pronounced ’Paddie Naboo’). This misinformation likely arose from the similarity between the characters’ names and appearances, as well as the occasional verbal slip-up by fans. As for the Galactic Senate’s sweetheart, Padwet Naboo’s name was actually that of her mother, who ran for office in her daughter’s name after Padwet’s untimely death. Padwet Naboo, who was actually the love interest of Senator Organa (R2-D2’s owner), was played by Padwet’s decoy, Padwet Decoy, who was actually a droid programmed to distract Jedi while the Jedi were busy on other matters. |

### v3

@@
 RAG adopts information retrieval techniques to inject external knowledge from semantic-relevant documents as input contexts.
-However, due to today’s Internet being flooded with numerous noisy and fabricating content, it is inevitable that RAG systems are vulnerable to these noises and prone to respond incorrectly.
-To this end, we propose to optimize the retrieval-augmented Generator with a Adversarial Tuning Multi-agent system (ATM).
-The ATM steers the Generator to have a robust perspective of useful documents for question answering with the help of an auxiliary Attacker agent.
-The Generator and the Attacker are tuned adversarially for several iterations.
+However, since today’s Internet is flooded with numerous noisy and fabricating content, it is inevitable that RAG systems are vulnerable to these noises and prone to respond incorrectly.
+To this end, we propose to optimize the retrieval-augmented Generator with an Adversarial Tuning Multi-agent system (ATM).
+The ATM steers the Generator to have a robust perspective of useful documents for question answering with the help of an auxiliary Attacker agent through adversarially tuning the agents for several iterations.
 After rounds of multi-agent iterative tuning, the Generator can eventually better discriminate useful documents amongst fabrications.
-The experimental results verify the effectiveness of ATM and we also observe that the Generator can achieve better performance compared to state-of-the-art baselines.
+The experimental results verify the effectiveness of ATM and we also observe that the Generator can achieve better performance compared to the state-of-the-art baselines.
+The code is available at https://github.com/chuhac/ATM-RAG.
 * ATM: Adversarial Tuning Multi-agent System Makes a Robust Retrieval-Augmented Generator
@@
 Lingyong Yan211footnotemark: 1 Haibo Shi2 Dawei Yin2 Lei Sha1††thanks: Corresponding author.
-1Beihang University, Beijing, China 2Baidu Inc., Beijing, China junda_zhu@outlook.com lingyongy@gmail.com shihaibo@baidu.com yindawei@acm.org shalei@buaa.edu.cn
+1Beihang University, Beijing, China 2Baidu Inc., Beijing, China junda_zhu@outlook.com lingyongy@gmail.com haiboshi@outlook.com yindawei@acm.org shalei@buaa.edu.cn
 1 Introduction
 Large Language Models (LLMs) such as Llama Touvron et al.
-(2023a, b), Mistral Jiang et al.
+(2023a, b); Dubey et al.
+(2024), Mistral Jiang et al.
 (2023), or GPT-4 Achiam et al.
 (2023) have demonstrated impressive power in massive artificial intelligence tasks, especially the question answering.
-However, due to the knowledge deficiency and dynamic updating, given knowledge-intensive or long-tail questions, LLMs often fail to provide concise answers, leading to rejection to answer Xu et al.
+However, due to knowledge deficiency and dynamic updating, given knowledge-intensive or long-tail questions, LLMs often fail to provide concise answers, leading to rejection to answer Xu et al.
 (2024) or hallucination Macpherson and Platchias (2013); Zhang et al.
@@
 (2020); Gao et al.
-(2021), and then injected into the prompts of LLMs (termed as the Generator) to generate final answers.
+(2021), and then injected into the context of LLMs (termed as the Generator) to generate final answers.
 Most retrieval-augmented Generators refer to multiple relevant documents in practice to ensure the comprehensiveness of final answers.
@@
 The main reason owes to a widely observed fact that most LLMs are vulnerable to noisy or fabricating knowledge Shi et al.
-(2023a); Cuconasu et al.
+(2023); Cuconasu et al.
 (2024); Wu et al.
@@
 We propose a multi-agent system and introduce an aggressive Attacker to improve the robustness of the Generator in RAG-QA.
-We propose a novel optimization method (termed MITO) to improve Generator’s robustness against LLM hallucination content and find it hopeful to improve the generation capacity and robustness of the Generator simultaneously.
+We propose a novel optimization method (termed MITO) to improve Generator’s robustness against LLM hallucination content and find it hopeful of improving the generation capacity and robustness of the Generator simultaneously.
 We further evaluate the Generator resisting various noises with comprehensive analysis, which is a strong endorsement of the validity of the proposed adversarial tuning and iterative optimization.
@@
 (2022) and self-revision, forming a reasoning chain towards generating the final answer.
-As a separate system, retriever and generator have different training objectives, giving rise to risks of indirect optimization.
+As a separate system, the retriever and generator have different training objectives, giving rise to risks of indirect optimization.
 To this end, REPLUG Shi et al.
-(2023b) tunes the retriever with the output token probability without the need for direct access to generator’s parameter.
+(2024) tunes the retriever with the output token probability without the need for direct access to generator parameters.
 RA-DIT Lin et al.
@@
 RAG-end2end Siriwardhana et al.
-(2023) proposes to dynamically re-index the document library with optimized retriever during training.
+(2023) proposes to dynamically re-index the document library with an optimized retriever during training.
 Self-RAG Asai et al.
@@
 (2024) trains models to be fluent in answer generation and aware of whether to retrieve and what to retrieve, mitigating noises brought by unnecessary retrieval.
-It is widely recognized that with more parameter, LLM also can act as a re-ranker better Sun et al.
-(2023).
+It is widely recognized LLM also can act as a re-ranker better Sun et al.
+(2023) with more parameters.
 Recent works like MetaEOL Lei et al.
@@
 (2024), LLM2Vec BehnamGhader et al.
-(2024) and NV-Embed Lee et al.
-(2024) train or prompt LLMs to act as a sentence encoder and discover that LLMs have strong embedding capabilities.
+(2024), NV-Embed Lee et al.
+(2024) and bge-en-icl Li et al.
+(2024) prompt, train or construct in-context learning examples for LLMs to act as a document encoder and showcase that LLMs have a strong embedding capability.
 REAR Wang et al.
-(2024a) further plugs an extra re-ranking module to LLMs, and simultaneously improve its document ranking and answer generation capabilities.
-2.2 Adversarial Learning and Robustness
+(2024a) further plugs an extra re-ranking module to LLMs, and simultaneously improves its document ranking and answer generation capabilities.
+2.2 Adversarial Learning and Robust RAG
 Generative Adversarial Networks Goodfellow et al.
@@
 Its setting makes it possible that the robustness of the discriminator can be gradually enhanced with the adversarial tuning proceeding.
-Similar idea also works in natural language processing.
+This idea also works in natural language processing.
 Li et al.
 (2017) utilizes the trajectory-level reward from the discriminator distinguishing machine-generated content and conduct REINFORCE Williams (1992) to enhance the generator’s anthropomorphism.
-Different from GAN-inspired methods mentioned above, we argue that, our work expands the adversarial idea to a system where both agents are generative language models.
 As for the RALMs, some recent works like RetRobust Yoran et al.
-(2023) are aimed at making LLMs robust against irrelevant documents.
-However, to the best of our knowledge, our proposed ATM is the first to consider the vulnerability of RAG systems to LLM fabricated knowledge and introduce and adversarial agent in the optimization of the Generator.
+(2024), RAAT Fang et al.
+(2024) and RobustRAG Xiang et al.
+(2024) are aimed at making LLMs robust against irrelevant and noisy documents.
+However, to the best of our knowledge, our proposed ATM is the first to consider the vulnerability of RAG systems to LLM fabricated content, which is prone to produce hallucinations.
+We introduce an adversarial agent in the optimization of the Generator on top of adversarial ideas mentioned above, and expand GAN-inspired methods to a system where both agents are generative language models.
 3 ATM System
@@
 3.1 Attacker
-With the primary goal of improving the robustness of the Generator to LLM-fabricating content, the Attacker should be able to inject fabrications into retrieved document list which can challenge the Generator successfully.
+With the primary goal of improving the robustness of the Generator to LLM-fabricating content, the Attacker should be able to inject fabrications into the retrieved document list which can challenge the Generator successfully.
 In addition, some studies Liu et al.
@@
 Fabrication Generation
-Provided queries and retrieved document lists, the Attacker generates semantically related but useless or incorrect fabrications iteratively, end up forming the attacked list containing originally retrieved documents and multiple fabricated documents.
+Provided queries and retrieved document lists, the Attacker generates semantically related but useless or incorrect fabrications iteratively, and ends up forming the attacked list containing originally retrieved documents and multiple fabricated documents.
 Concretely, the fabrication generation consists of multiple iterations.
@@
 the LLM in RAG) takes the user query together with retrieved or attacked document list as the inputs, and is aimed at remaining robust to the noises and generating correct answers.
-To this end, a robust Generator should be capable of identifying and leveraging all relevant and useful documents for generating and ignoring those noisy knowledge, regardless of given original or fabricating document list.
+To this end, a robust Generator should be capable of identifying and leveraging all relevant and useful documents for generating and ignoring noisy knowledge, regardless of given original or fabricating document list.
 In other words, the goal of the Generator can be formalized to maximize the following objective:
@@
 (2018) of the Generator, which is the token-level conditional probability of answer with token length as formalized in Equation 3 in detail; The is a function measuring the distance between two generation probabilities.
-can be further calculated by Equation.
+can be further calculated by Eqn.
 3:
@@
 Specifically, we fine-tune the Generator using annotated SFT data whose loss function is as follows:
-Apart from vanilla fine-tuning, we also perform three strategies to synthesize more training data based on original SFT samples: (1) answering original question with only one document, (2) answering without document, (3) ground-truth document extraction222The prompt templates can be found in Appendix A.2..
+Apart from vanilla fine-tuning, we also perform three strategies to synthesize more training data based on original SFT samples: (1) answering the original question with only one document, (2) answering without document, (3) ground-truth document extraction222The prompt templates can be found in Appendix A.2..
 As for the Attacker, we directly prompt it without initial tuning to generate the fabrications as described in Section 3.1.
@@
 where is similar to Equation 4 but uses the attacked document list as input.
-is the token-level Kullback–Leibler Divergence Kullback and Leibler (1951) of answer generating probabilities between given normal document list and attacked document list.
+is the token-level Kullback–Leibler Divergence Kullback and Leibler (1951) of answer generating probabilities between the given normal document list and the attacked document list.
 is a pre-set hyper-parameter.
-Math derivations can be found in Appendix B, implementation details can be found in Appendix D.2.
+Math derivations can be found in Appendix B, and implementation details can be found in Appendix D.2.
 5 Experiments
@@
 Self-RAG | |||||||||
-RetRobust | |||||||||
+RetRobust | 49.47 | ||||||||
+RAAT | 60.64 | ||||||||
 ATM | |||||||||
 ATM | |||||||||
-ATM | |||||||||
-ATM |
+ATM | 57.73 | 54.21 | 63.23 | 69.33 | 72.30 | 56.31 | |||
+ATM | 47.35 |
 5.1 Experimental Setup
@@
 (2024); Wang et al.
-(2024a) uses different settings (e.g., retriever, knowledge base, number of documents included) when assessing RAG systems and there is no unified benchmark to evaluate both the generation capacity and robustness.
+(2024a) uses different settings (e.g., retriever, knowledge base, number of documents included) when assessing RAG systems, there is no unified benchmark to evaluate both the generation capacity and robustness.
 Inspired by these studies, we conduct a novel benchmark considering both the generation capacity and robustness.
@@
 Details can be found at Appendix D.1..
-For the test set, we use the queries from the test splits of all four datasets, where PopQA is an unseen dataset for model assessment.
+For the test set, we use the queries from the test splits of all four datasets, where PopQA is an unseen dataset during training, for model assessment.
 Different from the training set, for each query, we first retrieve top-ranked documents from Wikipedia and construct some fabricated documents using powerful LLMs.
@@
 Baselines
-We compare our method with two state-of-the-art RALMs: REAR Wang et al.
-(2024a) which follows a rank-then-generate setting and Self-RAG Asai et al.
-(2023) which makes LLMs dynamically retrieve and generate answers.
-We also select a recent work–RetRobust Yoran et al.
-(2023), which is also aimed at improving LLMs’ robustness to irrelevant documents444We notice some contemporaneous work, i.e.
-RAAT Fang et al.
-(2024) and RobustRAG Xiang et al.
-(2024).
-Since they don’t have publicly accessible code or model weights, we didn’t reproduce their results in our experiments..
-Besides, we compare our model’s performance with Llama2 model without tuning to illustrate the improvement brought by the proposed optimization.
+We compare our method with four state-of-the-art RALMs: 1) REAR Wang et al.
+(2024a) which follows a rank-then-generate setting; 2) Self-RAG Asai et al.
+(2024) which makes LLMs self-perceptively retrieve external knowledge and generate answers; 3) RetRobust Yoran et al.
+(2024) which is aimed at improving LLMs’ robustness to irrelevant documents; 4) RAAT Fang et al.
+(2024) which enhances LLMs’ performance to generate answers and discriminate noisy documents through dual-task learning on the constructed dataset.
 LLMs | Natural Questions | TriviaQA | WebQuestions | ||||||
@@
 Self-RAG | |||||||||
-RetRobust | |||||||||
+RetRobust | 58.56 | 63.72 | |||||||
 ATM | |||||||||
-ATM | |||||||||
-ATM | |||||||||
-ATM | |||||||||
+ATM | 47.93 | 46.06 | 54.37 | ||||||
+ATM | 70.41 | 61.65 | 73.39 | ||||||
+ATM | 54.27 | ||||||||
 Llama3-B | |||||||||
@@
 Self-RAG | |||||||||
-RetRobust | |||||||||
+RetRobust | 52.29 | 49.14 | 58.17 | ||||||
 ATM | |||||||||
-ATM | |||||||||
-ATM | |||||||||
-ATM |
+ATM | 46.36 | 44.72 | 53.67 | ||||||
+ATM | 61.23 | ||||||||
+ATM | 52.69 | 64.40 |
 LLMs | PopQA | ||
@@
 ATM | |||
-ATM | |||
+ATM | 51.42 | 42.06 | 54.05 |
 ATM | |||
@@
 Take the ATM as an example, it achieves at least Subspan EM score improvements, EM score improvements and F1 score improvements on Natural Questions dataset.
-We can also observe similar tendencies on the other two datasets.
-It verifies that through the two-stage tuning, ATM Generators can achieve better performance facing noisy retrieval documents for RAG-QA.
+We can also observe similar tendencies on the other datasets.
+It verifies that through the two-stage tuning, ATM Generators can achieve better performance when facing noisy retrieval documents for RAG-QA.
 By comparing robustness-aware tuned models (i.e.
-ATM and RetRobust) with the trivially tuned RALMs (i.e.
-REAR and Self-RAG), we can observe considerable advancements on almost three datasets, which reveals that RALMs can benefit from robustness-aware learning to enhance its generation capacity.
-And ATM shows significant improvements compared with RetRobust, which serves as a strong endorsement of the effectiveness of the proposed adversarial tuning method.
+ATM, RetRobust and RAAT) with the trivially tuned RALMs (i.e.
+REAR and Self-RAG), we can observe considerable advancements on almost three datasets, which reveals that RALMs can benefit from robustness-aware learning to enhance their generation capacity.
+And ATM usually achieves significant improvements compared with RetRobust and RAAT, which serves as a strong endorsement of the effectiveness of the proposed adversarial tuning method.
 Besides, we find an obvious performance gap of REAR compared to its original paper encountering fabrications.
@@
 This also confirms the necessity for the robustness-aware optimization of the RALMs.
-In addition, we observe that the performance increases with the ATM optimization proceeding and eventually reach the convergence after at most three iterations.
+In addition, we can also observe that the performance increases with the ATM optimization proceeding and eventually reach the convergence after at most iterations.
 It is also noteworthy that Generator without adversarial tuning (i.e., ATM) can still achieve better or comparable performance.
@@
 ATM | |||||||||
+ATM | 69.41 | ||||||||
 ATM | |||||||||
-ATM | |||||||||
-ATM | |||||||||
-ATM | |||||||||
+ATM | 57.73 | 54.21 | 63.23 | 59.77 | 72.30 | 48.91 | 56.31 | ||
+ATM | 47.35 | ||||||||
 ATM | |||||||||
@@
 Full Attacker | ||
-w/o LP | - | |
-w/o FG | - |
+w/o LP | ||
+w/o FG |
 5.3 Detailed Analysis
-Robustness against Different Fabrication Generators
+Robustness against different fabrication generators.
 We further evaluate the robustness of RALMs against fabrications generated by different LLMs (i.e., Mixtral-B, Mixtral-B and Llama3-B) as reported in Table 1 and Table 2.
@@
 Although RetRobust performs better on Natural Questions when Llama3-B acts as the fabrication generator (ATM achieves comparable performance), it is hard to scale to other settings.
-Robustness against Different Fabrication Numbers
-Figure 4 illustrates the robustness of different models against different fabrication numbers in ten retrieved documents (varying from to ).
-As is illustrated, the ATM Generator can still surpass its baselines against different fabrication numbers.
-Our model performs more stably (with smoother curves) when the number of fabrications increase, especially the ATM and ATM.
-Performance on Unseen Dataset
-To alleviate the dataset bias, we conduct experiments that the RALMs are tested with unseen dataset PopQA.
+Robustness against different fabrication numbers.
+Figure 4 illustrates the robustness of different models against different fabrication numbers (varying from to ) in the case of a total of documents.
+As is illustrated, the ATM Generator has more stable performance (with smoother curves) when the number of fabrications increase, especially the ATM and ATM.
+When there is less fabrications, our model also performs comparably with state-of-the-art RALMs.
+And with the number of fabrications increasing, our model can surpass its baselines, showcasing its stability and robustness.
+Performance on unseen dataset.
+To alleviate the dataset bias, we conduct experiments to evaluate the RALMs on an unseen dataset–PopQA.
 The experimental results are shown in Table 3.
 And we can find that our methods can still surpass the baselines.
-Surprisingly, Self-RAG achieves the worst performance over EM and F1 scores, we induce that this may due to its self-reflection introducing extra noise of generations.
-Ablation Study
-We also conduct ablation to prove the necessity of different attacking types of the Attacker.
+Surprisingly, Self-RAG achieves the worst performance over EM and F1 scores, we induce that this may be due to its self-reflection introducing extra noise of generations.
+Visualization of attacking intensity.
+We also monitored the attacking intensity of Attacker as optimization proceeds.
+For the visualization analysis, we analyze the Log Loss of Generator given Attacker fabrications at different iterations.
+As is showcased in Figure 5, the increasing PPL of the Generator shows that Attacker is emerging increasingly stronger attacking patterns which greatly obstruct the Generator from generating right answers.
+In order to observe the change of attacking patterns more intuitively, we also conduct case study as reported in Appendix F.
+Ablation study.
+We also conduct ablation to verify the necessity of different attacking types of the Attacker.
 As shown in Table 5, the performance drops obviously without fabrication generation (FG), which injects noisy fabrications.
-And the list permutation (LP) is also proven to be necessary.
-We induce that the positional change can increase the diversity of document list, thus enhancing the training effectiveness.
-We also monitored the attacking intensity of Attacker across iterations as reported in Appendix F, including a visualization analysis and the case study.
-Influence of Hyper-parameter
+The list permutation (LP) is also proven to be necessary.
+We induce that the positional change can increase the diversity of the document list, thus enhancing the training effectiveness.
+Influence of hyper-parameters.
 We conduct experiments to investigate the influences of hyper-parameter settings.
@@
 As is formalized, when the optimization degenerates to SFT.
-From Table 4 we observe that, (1) with pure SFT the optimization has relatively low performance ceilings: with a drop observed on Natural Questions; (2) lower makes the optimization more steady while a higher one brings instability.
+From Table 4 we observe that, (1) with vanilla SFT the optimization has relatively low-performance ceilings: with a drop observed on Natural Questions; (2) lower makes the optimization more steady while a higher one brings instability.
 When becomes , a significant drop is witnessed at the start of optimization.
@@
 Though we encourage ATM Attacker to fabricate misleading knowledge, it is exactly what we seek to do in this work to mitigate the impact of retrieved fake LLM-generated content, which is believed to be particularly important in today’s RAG-QA systems.
-Acknowledgements
-This work was supported by the National Natural Science Foundation of China under grant No. KZ37117501, No. ZG216S23E8, No. 62306024, and No. 92367204.
+Acknowledgement
+This work was supported by the National Science Fund for Excellent Young Scholars (Overseas) under grant No. KZ37117501, National Natural Science Foundation of China (No. 62306024, No. 92367204).
 Appendix A Prompts of Agents
@@
 We utilize Contriever Gao et al.
-(2023) 555https://huggingface.co/facebook/contriever-msmarco further fine-tuned with MS MARCO Bajaj et al.
+(2023) 444https://huggingface.co/facebook/contriever-msmarco further fine-tuned with MS MARCO Bajaj et al.
 (2016) for passage retrieval.
 21M Wikipedia passages dumped on Dec.
-20, 2018 are adopted as our external knowledge base 666https://dl.fbaipublicfiles.com/dpr/wikipedia_split/psgs_w100.tsv.gz across datasets as listed in Section 5.
+20, 2018 are adopted as our external knowledge base 555https://dl.fbaipublicfiles.com/dpr/wikipedia_split/psgs_w100.tsv.gz across datasets as listed in Section 5.
 In order to accelerate the vector base similarity search, we utilize faiss Johnson et al.
@@
 To this end, we use mixed-precision training with a model in bfloat16 data type Kalamkar et al.
-(2019) implemented in apex777https://github.com/NVIDIA/apex and ZeRO Rajbhandari et al.
-(2020) stage 1 implemented with DeepSpeed888https://github.com/microsoft/DeepSpeed which splits optimizer state across GPUs thus saving memory.
+(2019) implemented in apex666https://github.com/NVIDIA/apex and ZeRO Rajbhandari et al.
+(2020) stage 1 implemented with DeepSpeed777https://github.com/microsoft/DeepSpeed which splits optimizer state across GPUs thus saving memory.
 Inference
-We utilize vLLM999https://github.com/vllm-project/vllm with optimized Paged Attention for LLM inference Kwon et al.
+We utilize vLLM888https://github.com/vllm-project/vllm with optimized Paged Attention for LLM inference Kwon et al.
 (2023) which is seamlessly compatible with the state-of-the-art LLM library transformers Wolf et al.
@@
 With more candidates injected, though the recall rate continuously improves, the accuracy of the Generator responses is easily bottlenecked by more noise.
-Appendix F Detailed Analysis of Attacker
-In order to have an intuition of the increasing attack intensity of Attacker, we conduct case study and visualization to analyze its attacking patterns.
-F.1 Visualization and Analysis
-For the visualization analysis, we analyze the Log Loss of Generator given Attacker fabrications across different iterations.
-As depicted in Figure 6, the increasing PPL of the Generator shows that Attacker is emerging increasingly stronger attacking patterns which greatly obstruct the Generator from generating right answers.
-F.2 Case Study
+Appendix F Case Study of Attacker
+In order to have an intuition of the increasing attack intensity of Attacker, we conduct case study to analyze its attacking patterns.
 As shown in Table 7 with fabrications provided by the Attacker across different optimization iterations.